x squared - 7x = -4
could you please explain how you would answer this
thank you

Answer by stanbon(75887)

x squared - 7x = -4
-----
x^2 -7x + 4 = 0
----
Use the Quadratic Formula to get:
x = [7 +- sqrt(49-4*1*4)]/2
-----
x = [7 +- sqrt(33)]/2
-----
x = (7/2) + (1/2)sqrt(33) or x = (7/2)-(1/2)sqrt(33)
===============
Cheers,
Stan H.